# Sim-to-real supervised domain adaptation for radioisotope identification

### v3

@@ Abstract
Spectroscopy and spectrometry are widely used characterization methods, and depending on the spectrum type, can illuminate details about a material’s crystallographic, molecular, elemental, or isotopic composition. The applications of such spectral characterization and analysis span fields including high-energy physics [1], remote sensing [2], biology [3], and nuclear material analysis [4]. While the motivations, physical mechanisms, and equipment used vary widely across spectral analysis, they are all unified by the processing of a “spectrum”–the histogram of discrete counts of a quantum arriving at a detector versus a physical quantity (usually energy or wavelength) indicative of each quantum over time. In the present work, we focus on gamma spectroscopy, which measures a spectrum of the number of gamma photons depositing energy in a detector across discrete ranges of deposited energy [5]. While we focus on specifics of gamma spectroscopy, the analysis of how to best enable spectral simulation-based inference is more general.
@@
Many detector types have been developed for detecting gammas and their energy, each with unique resolution (a measurement of the ability to discriminate between small differences in gamma energy), efficiency (a measurement of the proportion of gammas incident on the detector that are detected), and calibration (a description of how the deposition energy is converted to an electrical signal and eventually measured). Traditional techniques for gamma spectral analysis search for peaks in the measured gamma energy spectrum that can then be mapped to specific decay energies of a radioisotope [6, 7, 8]. These methods typically include numerous time-consuming preprocessing steps and require close monitoring by an expert spectroscopist [9, 10, 11]. Furthermore, classical approaches struggle in the presence of thick shielding or when the spectrum contains many overlapping material signatures. Supervised machine learning is an excellent alternative candidate for spectral analysis due to its ability to learn near-optimal representations of data for a given task [12]. Gamma spectroscopy, however, introduces some challenges to the adoption of supervised machine learning. First and foremost, supervised machine learning requires large labeled datasets for high performance, and so-called “neural scaling laws” have emerged showing enormous scale is required for difficult tasks (e.g., 38 million images per class for 99% accurate image classification on natural images [13]). The collection of such large datasets of gamma spectroscopy data would be extremely costly and time-consuming. Further, the labeling of a gamma spectrum is ambiguous. At the spectrum level, if a complete spectrum is labeled according to a recorded set of experimenter-placed radioactive sources, these labels will be incomplete due to naturally occurring radioactive material (NORM) in the background and cosmic radiation sources; at the quanta level, it is presently impossible to label each detection event according to its origin source from among the nearby radioactive sources, NORM, and cosmic sources [14].
@@
This has led to a trend in the field of supervised learning for spectroscopic analysis where simulation-based inference is leveraged [14, 15, 16, 17]. Amongst simulation-based inference techniques, naive application to real data of a classifier trained on simulated data incurs a huge misclassification risk, since the ability of such models to generalize is inherently limited by the difference in the data distributions between those generated by simulation and those generated by real processes, hereafter the “sim-to-real gap”. In this analysis, we consider the scenario where we have access to a data-rich source domain (simulations) and wish to transfer knowledge to perform classification in a data-scarce target domain (experimental measurements). In the present work, we characterize the effectiveness of this approach, leaving the investigation of other simulation-based inference approaches, such as the iterative approaches from Ref. [15] or unsupervised and semi-supervised approaches [18] for future studies. As such, we consider two primary research questions:
@@ 2 Related Work
Machine learning methods were first used for radioisotope identification by Olmos et al. [19, 20], where the authors demonstrated that artificial neural networks (ANNs) can achieve high classification accuracy on experimental gamma spectra. More recently, Liang et al. trained a convolutional neural network (CNN) using simulated mixed sources, achieving high accuracy on low-resolution spectra [21]. Kamuda developed an open-source Python package annsa, capable of creating simulated gamma-ray training datasets and applying machine learning models to solve spectroscopic tasks [22]. The author built upon this research in Ref. [23], finding that CNNs had smaller variance in their output than MLPs.
@@
Moore et al. applied the CNN architecture to time-sequenced gamma spectra by analyzing a two-dimensional waterfall plot [24]. The authors continued this research in Ref. [25], where they achieved mixed results leveraging transfer learning techniques to combine simulated data with experimental measurements. Daniel et al. also considered the task of out-of-distribution (OOD) testing, achieving accuracy on real data after training a CNN on a fully synthetic database [26].
@@
Bilton et al. examined the use of ANNs for mobile detection using simulated NaI(Tl)detectors, finding that autoencoder-based spectral anomaly detection outperforms PCA and NMF by up to and recurrent neural network enhanced source identification boosts accuracy by over static ANNs [27]. Ghawaly et al. similarly trained a deep convolutional autoencoder, outperforming principal component analysis (PCA) based methods at flagging anomalous gamma spectra in both real-world and synthetic urban environments [28].
@@
Khatiwada et al. surveyed different machine learning algorithms for emergency response applications. The authors found that models trained on synthetic datasets offered excellent performance on simulated test sets, and performance was comparable to conventional methods when applied to an experimental dataset [29]. Li et al. introduced a more advanced transformer-based neural network for nuclide identification, achieving an improved recognition rate compared to classical architectures [30]. The authors reshaped 1024-channel gamma spectra into patches, and the z-score-normalized counts of each patch were used as the embedding vector for the subsequent attention layers.
@@
Morrow et al. developed an open-source Python package PyRIID for generating synthetic gamma spectra for training machine learning models [31]. Van Omen et al. used PyRIID to synthesize short-lived fission products to train a model using a semi-supervised loss function, aiming to perform OOD multi-label proportion estimation [32]. Barradas et al. similarly used high-quality synthetic spectra to train a complex ANN, outperforming conventional peak-recognition methods when identifying up to ten radionuclides in multi-source LaBr3 gamma spectra [33].
@@ 3 Dataset Curation
We consider six datasets, divided into three domain adaptation scenarios, summarized in Table 1. In the first scenario (sim-to-sim adaptation), the task is multi-label proportion estimation on a simulated dataset of mixed geant spectra using a HPGe detector. In the second and third scenarios (sim-to-real adaptation), the task is multi-class, single-label classification on a dataset of experimental handheld spectra taken using LaBr3 and NaI(Tl) detectors.
@@
| Target | geant | Mixed | 55 | ||||||||
@@ 3.1 Sim-to-sim adaptation
In the first scenario, we simulated a dataset of gamma spectra using gadras (source domain) and aim to perform isotopic identification on a spectral dataset simulated in geant (target domain). In geant, we simulated the gamma emission spectra of 55 individual radioisotopes by uniformly dissolving each isotope in a scintillation cocktail. A high-purity germanium (HPGe) detector was placed one meter from the source, and the energy of all incoming gamma rays was recorded and binned to create a gamma spectrum. The simulations were repeated using a mix of different scintillation cocktails and HPGe detector models in order to capture additional variance which may be present in a realistic environment. We leveraged the G4ARES simulation framework to improve estimations of gamma cascade summing corrections [16]. In post-processing, we applied a low-energy cutoff and performed Gaussian energy broadening over each spectrum [5]. Furthermore, we introduced slight variations to energy calibration and detector resolution between different detector setups. Once these template spectra were simulated, we created mixed spectra by arbitrarily summing up to 14 individual spectra with random proportions. Our choice to include multiple isotopes per spectrum captures the common scenario in which several radioisotopes are simultaneously present in a material sample. We subsequently added background (cosmic, 40K, 226Ra, and 232Th) and Poisson noise. We synthesized a total of training spectra, along with separate holdout sets for validation and testing. To clarify, we did not use the entire dataset for model training; instead, we trained on data subsets of varying sizes to compare the performance of different models (with and without pretraining) as a function of dataset size.
We followed a similar sequence of steps for generating gadras data, first simulating 55 template spectra using a 95% efficiency HPGe detector. The spectra were then mixed, background-added, and Poisson resampled to produce a total of training spectra, performed using the PyRIID SeedMixer and StaticSynthesizer classes [31]. In both gadras and geant, all spectra consisted of 1024 uniform energy bins spanning from 0 to 3000 keV, with slight endpoint fluctuations due to variations in energy calibration.
@@ 3.2 Sim-to-real adaptation
In the experimental domain adaptation scenarios, we again use gadras as our source domain, but seek to perform isotopic classification using experimental spectra. The experimental dataset was collected in 2018 at Pacific Northwest National Laboratory to characterize the performance of handheld LaBr3 and NaI(Tl) detectors. 32 different source isotopes were used, with either steel, lead, polyethylene, or no shielding. The data were preprocessed to remove spectra that were corrupted, empty, or missing metadata, as well as to interpolate all spectra onto a uniform 1024-bin energy grid from 0 to 3000 keV. This resulted in a labeled dataset of 15,091 LaBr3 spectra and 10,440 NaI(Tl) spectra, partitioned into training, validation, and test sets with a 70/15/15 split. As before, we only considered subsets of the training data at a time to benchmark performance as a function of training size. It is worth mentioning that the experimental datasets available for this study are relatively simple, and 100% testing accuracy can be achieved while only training on training examples due to the nature of the data collection process (single-label, high signal-to-noise ratio). While a more complicated experimental dataset is desirable, we still expect the results of this research to apply broadly and refer to the sim-to-sim adaptation as a case study using complex, mixed sources.
The gadras dataset was synthesized in an attempt to best replicate the experimental configuration. Spectra were simulated for each of the 32 radioisotopes over a range of source-to-detector distances, detector heights, energy calibrations, and shielding configurations. These spectra were then mixed with gadras-simulated background (cosmic, 40K, 226Ra, and 232Th) to create seed templates using the PyRIID SeedMixer class. The mixing fractions were drawn from a Dirichlet distribution whose parameters were estimated by fitting measured background spectra as a non-negative combination of the four simulated component spectra. These mixed spectrum seed templates were then used to create a synthetic dataset using the PyRIID StaticSynthesizer class, varying over parameters such as background count rate, live time, and signal-to-noise ratio, with the values of each parameter approximated empirically. In total, we synthesized training spectra, validation spectra, and testing spectra in less than 15 minutes on a personal laptop. This process was repeated for each detector type (LaBr3, NaI(Tl)).
@@ 4.1 Model Architectures
We implemented four machine learning architectures for this analysis. Multilayer perceptrons (MLPs) and convolutional neural networks (CNNs) are commonly used in the gamma-spectroscopy literature and serve as simple, effective benchmarks. We also consider attention-based models (first introduced for neural machine translation [34]) and include a prior transformer-based neural network for nuclide identification, TBNN (Li) [30]. Finally, we introduce TBNN (ours), a more general transformer that modifies the Vision Transformer (ViT) [35] for source identification in gamma spectroscopy (Fig. 1). Below, we motivate each architecture class by the assumptions it encodes about spectral data and the implications for generalization:
@@
For data representable as sequences, many different sequence models have been proposed and tested, progressing from recurrent neural networks (RNNs) [12], Transformers [34], and recent selective state space models (SSMs) [36]. Transformers are the most common choice due to their ability to train in a fully parallel manner, and the amount of literature available characterizing them. In gamma spectroscopy, their attention mechanism should be able to determine higher order interactions between the number of counts in different channels, which is analogous to what gamma spectroscopists perform implicitly. Limitations in interaction range are often cited for Transformers, so widely separated peaks from a specific isotope may be difficult to capture.
In TBNN (Li) [30], the authors first reshape a 1024-dimensional spectral input into a sequence and then add fixed sinusoidal positional encodings. The resulting embeddings are then fed into a stack of multi-head self-attention blocks, with a final dense layer mapping to the predicted isotopes. We introduce a more general transformer-based neural network, TBNN (ours), featuring several modifications over Li et al. [30]:
@@ 4.2 Training Protocols
For each domain adaptation scenario, we trained three classes of models using the datasets generated in Section 3. The first class, source-only, was a single model simply trained on the entire source domain (gadras) dataset. The second class of models, target-only, were trained solely on the target-domain dataset (geant in the sim-to-sim scenario; experiment in the sim-to-real scenario). Within this class, we trained multiple different models, each using a random subset of the target domain data of size for , where is the total number of training examples. The third class of models, domain-adapted, were pretrained using the source-domain dataset, and subsequently fine-tuned using the target-domain data. As before, multiple domain-adapted models were trained with data subsets of varying sizes to evaluate the effect of dataset size on model performance. For all models, input spectra were first z-score normalized, which is a common data preprocessing technique in machine learning to stabilize training [37].

### v4

@@ Abstract
Spectroscopy and spectrometry are widely used characterization methods, and depending on the spectrum type, can illuminate details about a material’s crystallographic, molecular, elemental, or isotopic composition. The applications of such spectral characterization and analysis span fields including high-energy physics [nonakaPileupCorrectionsHigherorder2020], remote sensing [gaoNovelUnsupervisedSegmentation2017], biology [fliriBiologicalSpectraAnalysis2005], and nuclear material analysis [Hagen2022]. While the motivations, physical mechanisms, and equipment used vary widely across spectral analysis, they are all unified by the processing of a “spectrum”–the histogram of discrete counts of a quantum arriving at a detector versus a physical quantity (usually energy or wavelength) indicative of each quantum over time. In the present work, we focus on gamma spectroscopy, which measures a spectrum of the number of gamma photons depositing energy in a detector across discrete ranges of deposited energy [knoll2010radiation]. While we focus on specifics of gamma spectroscopy, the analysis of how to best enable spectral simulation-based inference is more general.
@@
Many detector types have been developed for detecting gammas and their energy, each with unique resolution (a measurement of the ability to discriminate between small differences in gamma energy), efficiency (a measurement of the proportion of gammas incident on the detector that are detected), and calibration (a description of how the deposition energy is converted to an electrical signal and eventually measured). Traditional techniques for gamma spectral analysis search for peaks in the measured gamma energy spectrum that can then be mapped to specific decay energies of a radioisotope [sampson2007plutonium, genie2000, peakeasy]. These methods typically include numerous time-consuming preprocessing steps and require close monitoring by an expert spectroscopist [elliot2020peakmap, darweesh2019study, croce2021improved]. Furthermore, classical approaches struggle in the presence of thick shielding or when the spectrum contains many overlapping material signatures. Supervised machine learning is an excellent alternative candidate for spectral analysis due to its ability to learn near-optimal representations of data for a given task [Goodfellow2016]. Gamma spectroscopy, however, introduces some challenges to the adoption of supervised machine learning. First and foremost, supervised machine learning requires large labeled datasets for high performance, and so-called “neural scaling laws” have emerged showing enormous scale is required for difficult tasks (e.g., 38 million images per class for 99% accurate image classification on natural images [hestnessDeepLearningScaling2017]). The collection of such large datasets of gamma spectroscopy data would be extremely costly and time-consuming. Further, the labeling of a gamma spectrum is ambiguous. At the spectrum level, if a complete spectrum is labeled according to a recorded set of experimenter-placed radioactive sources, these labels will be incomplete due to naturally occurring radioactive material (NORM) in the background and cosmic radiation sources; at the quanta level, it is presently impossible to label each detection event according to its origin source from among the nearby radioactive sources, NORM, and cosmic sources [fernandezImprovementNuclideDetection2025].
@@
This has led to a trend in the field of supervised learning for spectroscopic analysis where simulation-based inference is leveraged [fernandezImprovementNuclideDetection2025, cranmerFrontierSimulationbasedInference2020, archambault2023g4ares, pierson2022alpha]. Amongst simulation-based inference techniques, naive application to real data of a classifier trained on simulated data incurs a huge misclassification risk, since the ability of such models to generalize is inherently limited by the difference in the data distributions between those generated by simulation and those generated by real processes, hereafter the “sim-to-real gap”. In this analysis, we consider the scenario where we have access to a data-rich source domain (simulations) and wish to transfer knowledge to perform classification in a data-scarce target domain (experimental measurements). In the present work, we characterize the effectiveness of this approach, leaving the investigation of other simulation-based inference approaches, such as the iterative approaches from Ref. [cranmerFrontierSimulationbasedInference2020] or unsupervised and semi-supervised approaches [tarvainen2018meanteachersbetterrole] for future studies. As such, we consider two primary research questions:
@@ 2 Related Work
Machine learning methods were first used for radioisotope identification by Olmos et al. [Olmos1991, Olmos1992], where the authors demonstrated that artificial neural networks (ANNs) can achieve high classification accuracy on experimental gamma spectra. More recently, Liang et al. trained a convolutional neural network (CNN) using simulated mixed sources, achieving high accuracy on low-resolution spectra [Liang2019rapid]. Kamuda developed an open-source Python package annsa, capable of creating simulated gamma-ray training datasets and applying machine learning models to solve spectroscopic tasks [kamuda2019automated]. The author built upon this research in Ref. [kamuda2020comparison], finding that CNNs had smaller variance in their output than multilayer perceptrons (MLPs).
@@
Moore et al. applied the CNN architecture to time-sequenced gamma spectra by analyzing a two-dimensional waterfall plot [Moore2019application]. The authors continued this research in Ref. [moore2020transfer], where they achieved mixed results leveraging transfer learning techniques to combine simulated data with experimental measurements. Daniel et al. also considered the task of out-of-distribution (OOD) testing, achieving accuracy on real data after training a CNN on a fully synthetic database [daniel2020automatic].
@@
Bilton et al. examined the use of ANNs for mobile detection using simulated NaI(Tl) detectors, finding that autoencoder-based spectral anomaly detection outperforms principal component analysis (PCA) and non-negative matrix factorization by up to and recurrent neural networks (RNNs) enhanced source identification accuracy by over static ANNs [bilton2021neural]. Ghawaly et al. similarly trained a deep convolutional autoencoder, outperforming PCA-based methods at flagging anomalous gamma spectra in both real-world and synthetic urban environments [ghawaly2022characterization].
@@
Khatiwada et al. surveyed different machine learning algorithms for emergency response applications. The authors found that models trained on synthetic datasets offered excellent performance on simulated test sets, and performance was comparable to conventional methods when applied to an experimental dataset [khatiwada2023machine]. Li et al. introduced a more advanced transformer-based neural network for nuclide identification, achieving an improved recognition rate compared to classical architectures [Li2024]. The authors reshaped 1024-channel gamma spectra into patches, and the z-score-normalized counts of each patch were used as the embedding vector for the subsequent attention layers.
@@
Morrow et al. developed an open-source Python package PyRIID for generating synthetic gamma spectra for training machine learning models [Morrow2021PyRIID]. Van Omen et al. used PyRIID to synthesize short-lived fission products to train a model using a semi-supervised loss function, aiming to perform OOD multi-label proportion estimation [vanomen2024multilabel]. Barradas et al. similarly used high-quality synthetic spectra to train a complex ANN, outperforming conventional peak-recognition methods when identifying up to ten radionuclides in multi-source LaBr3 gamma spectra [Barradas2025].
@@ 3 Dataset Curation
We consider six datasets, divided into three domain adaptation scenarios, summarized in Table 1. In the first scenario (sim-to-sim adaptation), the task is multi-label proportion estimation on a simulated dataset of mixed Geant4 spectra using a HPGe detector. In the second and third scenarios (sim-to-real adaptation), the task is multi-class, single-label classification on a dataset of experimental handheld spectra taken using LaBr3 and NaI(Tl) detectors.
@@
| Target | Geant4 | Mixed | 55 | ||||||||
@@ 3.1 Sim-to-sim adaptation
In the first scenario, we simulated a dataset of gamma spectra using GADRAS (source domain) and aimed to perform isotopic identification on a spectral dataset simulated in Geant4 (target domain). GADRAS (Gamma Detector Response and Analysis Software [horne2016gadras]) is a fast, deterministic, semi-empirical package for analyzing gamma-ray detector responses to user-specific sources. Geant4 (GEometry ANd Tracking [allison2016recent]) is a general-purpose, open-source Monte Carlo toolkit for simulating the passage of particles through matter, capable of modeling detector responses in complex 3D geometries.
In Geant4, we simulated the gamma emission spectra of 55 individual radioisotopes by uniformly dissolving each isotope in a scintillation cocktail. This source configuration mirrors the experimental arrangement reported in Ref. [pierson2022alpha] and was selected to capture realistic self-attenuation and geometric effects. A high-purity germanium (HPGe) detector was placed one meter from the source, and the energy of all incoming gamma rays was recorded and binned to create a gamma spectrum. We leveraged G4ARES (Geant4 Advanced Radio-Emission Simulation Framework [archambault2023g4ares]) to improve estimations of gamma cascade summing corrections. In post-processing, we applied a low-energy cutoff and performed Gaussian energy broadening over each spectrum [knoll2010radiation]. Furthermore, we introduced slight variations to energy calibration and detector resolution between different detector setups. Once these template spectra were simulated, we created mixed spectra by arbitrarily summing up to 14 individual spectra with random proportions. Our choice to include multiple isotopes per spectrum captures the common scenario in which several radioisotopes are simultaneously present in a material sample. We subsequently added background (cosmic, 40K, 226Ra, and 232Th) and Poisson noise. We synthesized a total of training spectra, along with separate holdout sets for validation and testing. To clarify, we did not use the entire dataset for model training; instead, we trained on data subsets of varying sizes to compare the performance of different models (with and without pretraining) as a function of dataset size.
We followed a similar sequence of steps for generating GADRAS data, first simulating 55 template spectra using a 95% efficiency HPGe detector (relative to a 33-inch NaI(Tl)). The spectra were then mixed, background-added, and Poisson resampled to produce a total of training spectra, performed using the PyRIID SeedMixer and StaticSynthesizer classes [Morrow2021PyRIID]. In both GADRAS and Geant4, all spectra consisted of 1024 uniform energy bins spanning from 0 to 3000 keV, with slight endpoint fluctuations due to variations in energy calibration.
@@ 3.2 Sim-to-real adaptation
In the experimental domain adaptation scenarios, we again used GADRAS as our source domain, but sought to perform isotopic classification using experimental spectra. The experimental dataset was collected in 2018 at Pacific Northwest National Laboratory to characterize the performance of handheld LaBr3 and NaI(Tl) detectors. Thirty-two different source isotopes were employed in standardized source-shield configurations (none, steel, lead, polyethylene). The data was parsed using SpecUtils, a library for opening and manipulating spectrum files [SpecUtils-github], and subsequently preprocessed to remove spectra that were corrupted, empty, or missing metadata. Furthermore, all spectra were interpolated onto a uniform 1024-bin energy grid from 0 to 3000 keV. This resulted in a labeled dataset of 15,091 LaBr3 spectra and 10,440 NaI(Tl) spectra, partitioned into training, validation, and test sets with a 70/15/15 split. As before, we only considered subsets of the training data at a time to benchmark performance as a function of training size.
The experimental datasets used in this study are independent of the sim-to-sim Geant4 simulations and reflect a different measurement scenario. Furthermore, we remark that the experimental datasets available for this study are relatively simple, and 100% testing accuracy can be achieved while only training on training examples due to the nature of the data collection process (single-label, high signal-to-noise ratio). While a more complicated experimental dataset is desirable, we still expect the results of this research to apply broadly and refer to the sim-to-sim adaptation as a case study using complex, mixed sources.
The GADRAS dataset was synthesized to replicate the experimental configuration by selecting the corresponding built‑in handheld LaBr3/NaI(Tl) detector model from the GADRAS detector library. Spectra were simulated for each of the 32 radioisotopes over a range of source-to-detector distances, detector heights, energy calibrations, and shielding configurations. These spectra were then mixed with GADRAS-simulated background (cosmic, 40K, 226Ra, and 232Th) to create seed templates using the PyRIID SeedMixer class. The mixing fractions were drawn from a Dirichlet distribution whose parameters were estimated by fitting measured background spectra as a non-negative combination of the four simulated component spectra. These mixed spectrum seed templates were then used to create a synthetic dataset using the PyRIID StaticSynthesizer class, varying over parameters such as background count rate, live time, and signal-to-noise ratio, with the values of each parameter approximated empirically. In total, we synthesized training spectra, validation spectra, and testing spectra in less than 15 minutes on a commodity laptop. This process was repeated for each detector type (LaBr3, NaI(Tl)).
@@ 4.1 Model Architectures
We implemented four machine learning architectures for this analysis. MLPs and CNNs are commonly used in the gamma-spectroscopy literature and serve as simple, effective benchmarks. We also consider attention-based models (first introduced for neural machine translation [Vaswani2017]) and include a prior transformer-based neural network for nuclide identification, TBNN (Li) [Li2024]. Finally, we introduce TBNN (ours), a more general transformer that modifies the Vision Transformer (ViT) [dosovitskiy2021image] for source identification in gamma spectroscopy (Fig. 1). Below, we motivate each architecture class by the assumptions it encodes about spectral data and the implications for generalization:
@@
For data representable as sequences, many different sequence models have been proposed and tested, progressing from RNNs [Goodfellow2016], Transformers [Vaswani2017], and recent selective state space models (SSMs) [guMambaLinearTimeSequence2023]. Transformers are the most common choice due to their ability to train in a fully parallel manner, and the amount of literature available characterizing them. In gamma spectroscopy, their attention mechanism should be able to determine higher order interactions between the number of counts in different channels, which is analogous to what gamma spectroscopists perform implicitly. Limitations in interaction range are often cited for Transformers, so widely separated peaks from a specific isotope may be difficult to capture.
In TBNN (Li) [Li2024], the authors first reshape a 1024-dimensional spectral input into a sequence and then add fixed sinusoidal positional encodings. The resulting embeddings are then fed into a stack of multi-head self-attention blocks, with a final dense layer mapping to the predicted isotopes. We introduce a more general transformer-based neural network, TBNN (ours), featuring several modifications over Li et al. [Li2024]:
@@ 4.2 Training Protocols
For each domain adaptation scenario, we trained three classes of models using the datasets generated in Section 3. The first class, source-only, was a single model simply trained on the entire source domain (GADRAS) dataset. The second class of models, target-only, were trained solely on the target-domain dataset (Geant4 in the sim-to-sim scenario; experiment in the sim-to-real scenario). Within this class, we trained multiple different models, each using a random subset of the target domain data of size for , where is the total number of training examples. The third class of models, domain-adapted, were pretrained using the source-domain dataset, and subsequently fine-tuned using the target-domain data. As before, multiple domain-adapted models were trained with data subsets of varying sizes to evaluate the effect of dataset size on model performance. For all models, input spectra were first z-score normalized, which is a common data preprocessing technique in machine learning to stabilize training [Goodfellow-et-al-2016].